# Towards Fair and Efficient Public Transportation: A Bus Stop Model

### v1

@@ 5 Conclusion
Our work suggests two natural avenues for further research. First, it remains open how to compute JR solutions (or even determine if they exist) for instances of $\alpha$-BSP where $\alpha\neq 0$. An even harder question is whether the core is always non-empty. This resembles the situation in approval-based committee voting, where the same question is famously open (see, e.g., Lackner and Skowron, 2023). Another promising direction is to extend our model to more complex topologies. For instance, one can consider the setting where the set of potential stops (and terminals) is a subset of $\mathbb{Q}^{2}$ or the vertex set of a (planar) graph. Further ahead, an important research challenge is to develop a richer framework that can be used to reason about fairness in more realistic models of public transport.
@@ Acknowledgements
Martin Bullinger and Edith Elkind were supported by the AI Programme of The Alan Turing Institute. Edith Elkind is also supported by an EPSRC grant EP/X038548/1.
@@ Appendix A Limitations of Solutions
We now strengthen Proposition 2.5 by showing that minimum-cost solutions do not even provide approximate JR. For 00-BSP, we strengthen this result even further by showing incompatibility between approximate JP and approximate cost-minimization. We start by formally defining our approximation objectives.
@@
Let $\alpha\in[0,1)$ and $\beta,\gamma\geq 1$. Then there exists an instance of $\alpha$-BSP such that every feasible solution minimizing total cost violates $\beta$-JR. Moreover, there exists an instance of 00-BSP such that no feasible solution simultaneously provides $\beta$-JR and $\gamma$-approximation of minimum cost.
@@
Let $\alpha\in[0,1)$. Consider the instance $\langle N,V,b,(\theta_{i})_{i\in N}\rangle$, depicted in Figure 6. The instance is defined based on two parameters $k,\ell\in\mathbb{N}$; the values of these parameters will be specified later. Let $N_{A}=\{a_{i}\colon i\in[k]\}$, $N_{D}=\{d_{i}\colon i\in[\ell]\}$, and set $N=N_{A}\cup N_{D}$. Let $V=\{-1,0\}\cup\{ik\colon i\in[k+1]\}$, and set $b=k+2$. Note that $b=|V|-1$. The agents’ types are as follows: For $i\in[k]$ we have $\theta_{a_{i}}=(0,ik)$ and $\theta_{d}=(-1,(k+1)k)$ for all $d\in N_{D}$.
@@
[…]
Figure 6: Incompatibility of approximate efficiency and approximate JR in Example 2.4 and Proposition 2.5.
@@
We are now ready to compute approximation guarantees. Consider target approximation ratios $\beta,\gamma\geq 1$, and set $\ell=\left\lceil 4\beta\right\rceil$ and $k=\left\lceil\gamma\ell\right\rceil+1$. Note that this implies $k>\ell$. Hence, by Equations 4 and 5, it holds that $c(S^{*})<c(S)$ for every feasible solution $S\neq S^{*}$, i.e., $S^{*}$ is the unique feasible solution of minimum cost. Moreover, if $\alpha=0$, it holds that $\frac{c(S)}{c(S^{*})}\geq\frac{k}{\ell}>\frac{\left\lceil\gamma\ell\right\rceil}{\ell}\geq\gamma$ for every feasible solution $S\neq S^{*}$. Thus, in this case $S^{*}$ is the unique feasible solution that is a $\gamma$-approximation of minimum cost.
@@
Appendix B Proofs Missing From Section 3
@@
Now, let $2\leq h\leq m$ and suppose that the claim is true for $h^{\prime}\in\{1,\dots,h-1\}$. First, by our initialization, the claim is correct for $c=0$. Consider $c\in\{1,\dots,b\}$, and let $h^{*}\in\{0,\dots,h-1$}. Then, $\mathbf{opt}[h,c]\leq\mathbf{opt}[h^{*},c-1]-\Delta[h^{*},h]$. Indeed, by starting with an optimal solution for $h^{*}$ and $c-1$ and adding a stop at $h$ we obtain a valid solution for $h$ and $c$. Hence, the cost of this solution is an upper bound on the minimum cost for $h$ and $c$. Consequently,
@@
Now, let $S$ be a solution of minimum cost where $h$ is the rightmost open stop and a budget of at most $c$ is used. Let $h^{*}$ be the second stop from the right in $S$ where we set $h^{*}=0$ if $S=\{h\}$. Then, $S^{*}:=S\setminus\{h\}$ is a candidate solution with budget $c-1$ where $h^{*}$ is the rightmost opened stop. So, $c(S^{*})\geq\mathbf{opt}[h^{*},c-1]$. Hence,
@@
[…]
@@
By the choice of the parameter $D$, the distance from $\ell_{i}$ to the closest stop in $V$ to its left is at least $1+\max_{j\in[k]}v_{j}$. Similarly, the distance from $r_{i}$ to the closest stop in $V$ to its right is at least $1+\max_{j\in[k]}v_{j}$. Hence, since $r_{i}-\ell_{i}=v_{i}$, an agent will take the bus if and only if both $\ell_{i}$ and $r_{i}$ are selected. We refer to this fact as Observation ()(*)( ∗ ).
@@
Conversely, assume that $\mathcal{I}$ admits a feasible solution $S\subseteq V$ of total cost at most $q$. Consider $I=\{i\in[k]\colon\{\ell_{i},r_{i}\}\subseteq S\}$. By Observation ()(*)( ∗ ), for each agent $i\in N\setminus I$ there is no faster way to commute from $\ell_{i}$ to $r_{i}$ than by walking. Hence, $c(S)=\sum_{i\in[k]}v_{i}-\sum_{i\in I}(1-\alpha)v_{i}$. Since $c(S)\leq q=\sum_{i\in[k]}v_{i}-(1-\alpha)v^{*}$, it follows that $\sum_{i\in I}v_{i}\geq v^{*}$. Moreover, $S$ is a feasible solution for $\mathcal{I}$. Hence, $\sum_{i\in I}w_{i}=\frac{1}{2}\sum_{i\in I}[\gamma(\ell_{i})+\gamma(r_{i})]\leq\frac{1}{2}\sum_{v\in S}\gamma(v)\leq\frac{1}{2}b=w^{*}$. Thus, $I$ certifies that $\mathcal{I}^{\text{Kn}}$ is a yes-instance of Knapsack. ∎
@@ C.1 Verification of Fair Solutions
    $S$ provides JR if and only if there does not exist a set $T\subseteq V$ with $|T|=2$ such that |$|\{i\in N\colon T\in\mathcal{P}_{i}(S)\}|\geq\frac{2n}{b}$.
@@
    $S$ is in the core if and only if there exists no set $T\subseteq V$ with $T\neq\varnothing$ such that |$|\{i\in N\colon\exists\,T^{\prime}\in\mathcal{P}_{i}(S)\text{ with }T^{\prime}\subseteq T\}|\geq\frac{|T|n}{b}$.
@@
[…]
@@
A solution $S$ is in the core if and only if its corresponding integer program (CoreTesting) has an optimal value of 00.
@@
Hence, a feasible solution of (CoreTesting) gives rise to a set of agents $M=\{i\in N\colon x_{i}=1\}$ and a set of stops $T=\{s\in V\colon y_{s}=1\}$ such that $M\subseteq\{i\in N\colon\exists\,T^{\prime}\in\mathcal{P}_{i}(S)\text{ with }T^{\prime}\subseteq T\}$ (by the first two constraints). If the optimal solution has value different from 00, then $T\neq\varnothing$ (otherwise, the first constraint cannot be satisfied) and |$|\{i\in N\colon\exists\,T^{\prime}\in\mathcal{P}_{i}(S)\text{ with }T^{\prime}\subseteq T\}|\geq\frac{|T|n}{b}$ (by the third constraint). By Proposition C.1, $S$ is not in the core.
Conversely, if $S$ is not in the core, we use Proposition C.1 to find $T\subseteq V$ with $T\neq\varnothing$ and |$|\{i\in N\colon\exists\,T^{\prime}\in\mathcal{P}_{i}(S)\text{ with }T^{\prime}\subseteq T\}|\geq\frac{|T|n}{b}$. We set $x_{i}=1$ if and only there exists $T^{\prime}\in\mathcal{P}_{i}(S)$ with $T^{\prime}\subseteq T$, $y_{s}=1$ if and only if $s\in T$, and $y_{T^{\prime}}=1$ if and only if $T^{\prime}\subseteq T$. This leads to a feasible solution of objective value greater than 00. Hence, the optimal solution has an objective value greater than 00. ∎
@@ C.2 Detailed Analysis of Frequency of Fairness Violations
Figure 8: Aggregated frequency of fairness violations of the solutions computed by Algorithm 1 and our benchmark algorithm. We aggregate the frequency for a fixed number of $5$, $10$, $15$, and $20$ agents (from top left to bottom right). The $x$-axis shows our range for the cost parameter $\alpha$ and the $y$-axis shows the percentage of the instances in which the desired property is not satisfied along with the standard error.
@@
Figure 10: Heat maps showing the frequency of core violations of the solutions computed by Algorithm 1 for pairs of $m$ and $b$ in instances of 00-BSP, $0.1$-BSP, $0.5$-BSP, and $0.9$-BSP (columns) and for JR and the core (rows). Each cell is averaged over all values of $n$.
Finally, we provide heat maps similar to the one in Figure 5 for different values of the cost parameter $\alpha$. Figure 9 and Figure 10 display the frequency of JR and core violations, respectively. For both fairness notions, we display the cases for $\alpha=0.1$, $\alpha=0.5$, and $\alpha=0.9$. For the core, the first picture is a version of Figure 5 where we adjusted the color scale to facilitate the comparison with other values of $\alpha$. Since Algorithm 1 always computes solutions that provide JR for 00-BSP, we omit a fourth picture for JR.
@@ C.3 Second Benchmark
(b) Aggregated frequency of fairness violations of the solutions computed by our proposed algorithm and the greedy algorithm. We aggregate the frequency for a fixed number of $5$, $10$, $15$, and $20$ agents (from top left to bottom right).
@@ Appendix D Approximately Stable Committee Selection
Definition D.2 (Stable Committees (or the Core) (Jiang et al., 2020)).
@@
Definition D.3 ($c$-Approximately Stable Committees (Jiang et al., 2020)).

### v2

@@ 5 Conclusion
Our work suggests two natural avenues for further research. First, it remains open how to compute JR solutions for instances of $\alpha$-BSP where $\alpha\neq 0$; indeed, we do not even know if such solutions always exist. An even harder question is whether the core is always non-empty. This resembles the situation in approval-based committee voting, where the same question is famously open (see, e.g., Lackner and Skowron, 2023). Another promising direction is to extend our model to more complex topologies. For instance, one can consider the setting where the set of potential stops (and terminals) is a subset of $\mathbb{Q}^{2}$ or the vertex set of a (planar) graph. Further ahead, an important research challenge is to develop a richer framework to reason about fairness in more realistic models of public transport.
@@ Acknowledgements
Most of this work was done while Edith Elkind was at University of Oxford. Martin Bullinger was supported by the AI Programme of The Alan Turing Institute. Edith Elkind and Mohamad Latifian were supported by the UK Engineering and Physical Sciences Research Council (EPSRC) grants EP/X038548/1 and EP/Y003624/1. We would like to thank the anonymous reviewers from AAMAS.
@@ Appendix A Limitations of Solutions
We now strengthen Proposition 2.5 by showing that minimum-cost solutions do not even provide approximate JR. For 0-BSP, we strengthen this result even further by showing incompatibility between approximate JP and approximate cost-minimization. We start by formally defining our approximation objectives.
@@
Let $\alpha\in[0,1)$ and $\beta,\gamma\geq 1$. Then there exists an instance of $\alpha$-BSP such that every feasible solution minimizing total cost violates $\beta$-JR. Moreover, there exists an instance of 0-BSP such that no feasible solution simultaneously provides $\beta$-JR and $\gamma$-approximation of minimum cost.
@@
Let $\alpha\in[0,1)$. Consider the instance $\langle N,V,b,(\theta_{i})_{i\in N}\rangle$, depicted in Figure 6. The instance is defined based on two parameters $k,\ell\in\mathbb{N}$; the values of these parameters will be specified later. Let $N_{A}=\{a_{i}\colon i\in[k]\}$, $N_{D}=\{d_{i}\colon i\in[\ell]\}$, and set $N=N_{A}\cup N_{D}$. Let $V=\{-1,0\}\cup\{ik\colon i\in[k+1]\}$, and set $b=k+2$. Note that b=|V|1b=|V|-1. The agents’ types are as follows: For $i\in[k]$ we have $\theta_{a_{i}}=(0,ik)$ and $\theta_{d}=(-1,(k+1)k)$ for all $d\in N_{D}$.
@@
We are now ready to compute approximation guarantees. Consider target approximation ratios $\beta,\gamma\geq 1$, and set $\ell=\left\lceil 4\beta\right\rceil$ and $k=\left\lceil\gamma\ell\right\rceil+1$. Note that this implies $k>\ell$. Hence, by Equations 5 and 6, it holds that $c(S^{*})<c(S)$ for every feasible solution $S\neq S^{*}$, i.e., $S^{*}$ is the unique feasible solution of minimum cost. Moreover, if $\alpha=0$, it holds that $\frac{c(S)}{c(S^{*})}\geq\frac{k}{\ell}>\frac{\left\lceil\gamma\ell\right\rceil}{\ell}\geq\gamma$ for every feasible solution $S\neq S^{*}$. Thus, in this case $S^{*}$ is the unique feasible solution that is a $\gamma$-approximation of minimum cost.
@@
[…]
Figure 6: Incompatibility of approximate efficiency and approximate JR in Theorem A.3.
@@
Appendix B Proofs Missing Concerning Efficiency in Section 3
@@
Now, let $2\leq h\leq m$ and suppose that the claim is true for $h^{\prime}\in\{1,\dots,h-1\}$. First, by our initialization, the claim is correct for $c=0$. Consider $c\in\{1,\dots,b\}$, and let $h^{*}\in\{0,\dots,h-1$}. Then, $\mathbf{opt}[h,c]\leq\mathbf{opt}[h^{*},c-1]-\Delta[h^{*},h]$. Indeed, by starting with an optimal solution for $h^{*}$ and c1c-1 and adding a stop at $h$ we obtain a valid solution for $h$ and $c$. Hence, the cost of this solution is an upper bound on the minimum cost for $h$ and $c$. Consequently,
@@
Now, let $S$ be a solution of minimum cost where $h$ is the rightmost open stop and a budget of at most $c$ is used. Let $h^{*}$ be the second stop from the right in $S$ where we set $h^{*}=0$ if $S=\{h\}$. Then, $S^{*}:=S\setminus\{h\}$ is a candidate solution with budget c1c-1 where $h^{*}$ is the rightmost opened stop. So, $c(S^{*})\geq\mathbf{opt}[h^{*},c-1]$. Hence,
@@
It is easy to see that our problem is in $\mathsf{NP}$: given a solution to $\alpha$-BSP, one can efficiently check if it satisfies the budget constraint and whether its cost does not exceed the cost bound.
For $\mathsf{NP}$-hardness, we provide a reduction from the $\mathsf{NP}$-complete problem Knapsack (Karp, 1972). An instance of Knapsack consists of a sequence of of $k$ weights $w=(w_{1},\dots,w_{k})\in\mathbb{N}^{k}$, $k$ values $v=(v_{1},\dots,v_{k})\in\mathbb{N}^{k}$, a capacity $w^{*}\in\mathbb{N}$, and a goal value $v^{*}\in\mathbb{N}$. An instance is a yes-instance if there exists a set of indices $I\subseteq[k]$ such that $\sum_{i\in I}v_{i}\geq v^{*}$ while $\sum_{i\in I}w_{i}\leq w^{*}$, and a no-instance otherwise.
@@
By the choice of the parameter $D$, the distance from $\ell_{i}$ to the closest stop in $V$ to its left is at least $1+\max_{j\in[k]}v_{j}$. Similarly, the distance from $r_{i}$ to the closest stop in $V$ to its right is at least $1+\max_{j\in[k]}v_{j}$. Hence, since $r_{i}-\ell_{i}=v_{i}$, an agent will take the bus if and only if both $\ell_{i}$ and $r_{i}$ are selected. We refer to this fact as Observation ()(*).
@@
Conversely, assume that $\mathcal{I}$ admits a feasible solution $S\subseteq V$ of total cost at most $q$. Consider $I=\{i\in[k]\colon\{\ell_{i},r_{i}\}\subseteq S\}$. By Observation ()(*), for each agent $i\in N\setminus I$ there is no faster way to commute from $\ell_{i}$ to $r_{i}$ than by walking. Hence, $c(S)=\sum_{i\in[k]}v_{i}-\sum_{i\in I}(1-\alpha)v_{i}$. Since $c(S)\leq q=\sum_{i\in[k]}v_{i}-(1-\alpha)v^{*}$, it follows that $\sum_{i\in I}v_{i}\geq v^{*}$. Moreover, $S$ is a feasible solution for $\mathcal{I}$. Hence, $\sum_{i\in I}w_{i}=\frac{1}{2}\sum_{i\in I}[\gamma(\ell_{i})+\gamma(r_{i})]\leq\frac{1}{2}\sum_{v\in S}\gamma(v)\leq\frac{1}{2}b=w^{*}$. Thus, $I$ certifies that $\mathcal{I}^{\text{Kn}}$ is a yes-instance of Knapsack. ∎
@@ C.1 Verification of Fair Solutions
    $S$ provides JR if and only if there does not exist a set $T\subseteq V$ with $|T|=2$ such that $|\{i\in N\colon T\in\mathcal{P}_{i}(S)\}|\geq\frac{2n}{b}$.
@@
    $S$ is in the core if and only if there exists no set $T\subseteq V$ with $T\neq\varnothing$ such that $|\{i\in N\colon\exists\,T^{\prime}\in\mathcal{P}_{i}(S)\text{ with }T^{\prime}\subseteq T\}|\geq\frac{|T|n}{b}$.
@@
$\text{s.t. }\quad x_{i}$ $\leq\sum_{T^{\prime}\in\mathcal{P}_{i}(S)}y_{T^{\prime}}$ $\forall\,i\in N$
@@
A solution $S$ is in the core if and only if its corresponding integer program (CoreTesting) has an optimal value of 0.
@@
Hence, a feasible solution of (CoreTesting) gives rise to a set of agents $M=\{i\in N\colon x_{i}=1\}$ and a set of stops $T=\{s\in V\colon y_{s}=1\}$ such that $M\subseteq\{i\in N\colon\exists\,T^{\prime}\in\mathcal{P}_{i}(S)\text{ with }T^{\prime}\subseteq T\}$ (by the first two constraints). If the optimal solution has value different from 0, then $T\neq\varnothing$ (otherwise, the first constraint cannot be satisfied) and $|\{i\in N\colon\exists\,T^{\prime}\in\mathcal{P}_{i}(S)\text{ with }T^{\prime}\subseteq T\}|\geq\frac{|T|n}{b}$ (by the third constraint). By Proposition C.1, $S$ is not in the core.
Conversely, if $S$ is not in the core, we use Proposition C.1 to find $T\subseteq V$ with $T\neq\varnothing$ and $|\{i\in N\colon\exists\,T^{\prime}\in\mathcal{P}_{i}(S)\text{ with }T^{\prime}\subseteq T\}|\geq\frac{|T|n}{b}$. We set $x_{i}=1$ if and only there exists $T^{\prime}\in\mathcal{P}_{i}(S)$ with $T^{\prime}\subseteq T$, $y_{s}=1$ if and only if $s\in T$, and $y_{T^{\prime}}=1$ if and only if $T^{\prime}\subseteq T$. This leads to a feasible solution of objective value greater than 0. Hence, the optimal solution has an objective value greater than 0. ∎
@@ C.2 Detailed Analysis of Frequency of Fairness Violations
Figure 8: Aggregated frequency of fairness violations of the solutions computed by Algorithm 1 and our benchmark algorithm. We aggregate the frequency for a fixed number of $5$, $10$, $15$, and 2020 agents (from top left to bottom right). The $x$-axis shows our range for the cost parameter $\alpha$ and the $y$-axis shows the percentage of the instances in which the desired property is not satisfied along with the standard error.
@@
Figure 10: Heat maps showing the frequency of core violations of the solutions computed by Algorithm 1 for pairs of $m$ and $b$ in instances of 0-BSP, $0.1$-BSP, $0.5$-BSP, and $0.9$-BSP (columns) and for JR and the core (rows). Each cell is averaged over all values of $n$.
Finally, we provide heat maps similar to the one in Figure 5 for different values of the cost parameter $\alpha$. Figure 9 and Figure 10 display the frequency of JR and core violations, respectively. For both fairness notions, we display the cases for $\alpha=0.1$, $\alpha=0.5$, and $\alpha=0.9$. For the core, the first picture is a version of Figure 5 where we adjusted the color scale to facilitate the comparison with other values of $\alpha$. Since Algorithm 1 always computes solutions that provide JR for 0-BSP, we omit a fourth picture for JR.
@@ C.3 Second Benchmark
(b) Aggregated frequency of fairness violations of the solutions computed by our proposed algorithm and the greedy algorithm. We aggregate the frequency for a fixed number of $5$, $10$, $15$, and 2020 agents (from top left to bottom right).
@@ Appendix D Approximately Stable Committee Selection
Definition D.2 (Jiang et al., 2020).
@@
Definition D.3 (Jiang et al., 2020).